# Optimizing Microgrid Composition for Sustainable Data Centers

### v1

@@ Abstract.
conference: Sustainable Supercomputing Workshop at SC25; Nov 16-21, 2025; St. Louis, MOcopyright: rightsretainedccs: Computing methodologies Simulation environmentsccs: Social and professional topics Sustainabilityfootnotetext: Preprint under review.
@@ 1. Introduction
On-site energy generation, in particular co-locating data centers with microgrids, offers a promising solution by aligning data center loads with local renewable energy resources, effectively reducing reliance on grid energy. While industry initiatives, like Google’s Partnership with Intersect Power and TPG Rise Climate to build renewable-driven data centers (Porat, 2024), demonstrate the feasibility of a commitment to co-located renewable energy generation, there is a lack of publicly available tools for modeling and planning such systems. Existing work has focused either on operational optimization or grid-connected strategies. For example, Carbon Explorer (Acun et al., 2023) introduces a holistic framework to balance embodied and operational carbon for 24/7 carbon-free data center operation. However, it relies on power purchase agreements and does not support automated right-sizing of on-site renewable microgrids. Other approaches, such as SHIELD (Qi et al., 2023), optimize resource allocation across geo-distributed data centers but do not address the integration or design of on-site energy systems. Similarly, Si et al. (Si et al., 2024) propose a blockchain-based control architecture for data center microgrids, focusing on real-time energy dispatch rather than infrastructure optimization.
@@
All code and data will be published after acceptance.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Workshops of the International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: Workshops of the International Conference for High Performance Computing, Networking, Storage and Analysis (SC Workshops ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3731599.3767562isbn: 979-8-4007-1871-7/2025/11ccs: Computing methodologies Simulation environmentsccs: Social and professional topics Sustainability
@@ 1. Introduction
On-site energy generation, in particular co-locating data centers with microgrids, offers a promising solution by aligning data center loads with local renewable energy resources, effectively reducing reliance on grid energy. While industry initiatives, like Google’s Partnership with Intersect Power and TPG Rise Climate to build renewable-driven data centers (Porat, 2024), demonstrate the feasibility of a commitment to co-located renewable energy generation, there is a lack of publicly available tools for modeling and planning such systems. Existing work has focused either on operational optimization or grid-connected strategies. For example, Carbon Explorer (Acun et al., 2023) introduces a framework to balance embodied and operational carbon for 24/7 carbon-free data center operation. However, it relies on power purchase agreements and does not support automated right-sizing of on-site renewable microgrids. Other approaches, such as SHIELD (Qi et al., 2023), optimize resource allocation across geo-distributed data centers but do not address the integration or design of on-site energy systems. Similarly, Si et al. (Si et al., 2024) propose a blockchain-based control architecture for data center microgrids, focusing on real-time energy dispatch rather than infrastructure optimization.
@@
All results reported in this paper are fully reproducible. The code and data are available at https://github.com/dos-group/vessim-opt.